OpenStudy (anonymous):

solve each system 2x-y+z =7 y+z =5

OpenStudy (anonymous):

Is this linear algebra?

OpenStudy (anonymous):

system of equations and inequalities

OpenStudy (anonymous):

Do you know what a matrix is?

OpenStudy (anonymous):

havt gotten there yet

OpenStudy (anonymous):

Your system of equations in this case is not complete. You have 3 variables but only 2 equations.

OpenStudy (anonymous):

This means you need to have a free variable.

OpenStudy (anonymous):

Are you familiar with this concept?

OpenStudy (anonymous):

that is the question to be answered and i think adding those two will also eliminate y from the equation

OpenStudy (anonymous):

i was supposed to be a three variable then u add the first two to get a different equation and u add the first and third to get another equation then u solve for variables

OpenStudy (anonymous):

not three variable but three equations

OpenStudy (anonymous):

You can't create a third equation from the existing equations, because that would be redundant.

OpenStudy (anonymous):

It's like saying: \[ x+y=2\\ 2x+2y=4 \]They're not independent equations, so they might as well be only a single equation.

OpenStudy (anonymous):

In any case, add the equations together to get: \[ 2x+2z=12 \implies x=6-z \]

OpenStudy (anonymous):

Next, subtract \(z\) from both sides of the second equation: \[ y=5-z \]

OpenStudy (anonymous):

Our solution to the system will be \((x,y,z) = (6-z,5-z,z)\). We need the free variable \(z\) because we just don't have enough equations to solve for it.